The wavelength of the green laser of the light effect device is 531 nm.

What is the frequency of such light?


1
Expert's answer
2021-08-13T10:21:01-0400

"λ = 531 \;nm = 531 \times 10^{-9} \;m \\

f=\frac{c}{λ}"

Light velocity "c = 3 \times 10^8 \;m/s"

"f = \frac{3 \times 10^8}{531 \times 10^{-9}} = 5.65 \times 10^{14}\;Hz"

Answer: "5.65 \times 10^{14}\;Hz"